Solve the equation by completing the square.
a. Write irrational numbers in simplest radical form.
b. Describe each step as you go.
X^2 + 4x – 8 = 0

Answer by stanbon(75887)

X^2 + 4x – 8 = 0
----
x^2 + 4x + ? = 8 + ?
----
x^2 + 4x + 4 = 8 + 4
-----
(x+2)^2 = 12
---
x+2 = +/-2sqrt(3)
------
x = -2+2sqrt(3) or x = -2-2sqrt(3)
=======================================
Cheers,
Stan H.
==============